If the world’s population is increasing at an annual rate of 1.5%, and there were 8 billion people in the year 2010, then in what year will the world’s population be 20 billion?
Answer #1

Change in population = 20 - 8 = 12 billion

Assuming exponential growth model

P = A*e^(r*t)

r = 1.5% = 0.015

A = 8 billion, P = 20 billion

20 = 8*e^(0.015*t)

2.5 = e^(0.015*t)

ln 2.5 = 0.015*t

t = (ln 2.5)/0.015

t = 61.09 years